Jump to main content.


Drake Chemical

EPA Daily Situation Reports for November 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Monday, November 30, 1998

FOR DATES: From 7:00 a.m. Sunday, Nov. 29, 1998 to 7:00 a.m. Monday, Nov. 30, 1998.

WEATHER: Overcast. The temperature range was 37oF to 62oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.45 hours and 759.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.8 tons per hour while on line. Downtime was due to flameout of SCC burners. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

Received ash analytical reports (bot-ash-11-23-98-288, bot-ash-11-24-98-289 and fly-ash-11-24-98-058). Bottom ash met project requirements for unrestricted backfill on-site. Fly ash analytical results indicated that the fly ash met the project specifications for restricted backfilling on-site due to the concentrations of chromium, cadmium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for David Modricker (USACE)


Sunday, November 29, 1998

FOR DATES: From 7:00 a.m. Saturday, Nov. 28, 1998 to 7:00 a.m. Sunday, Nov. 29, 1998.

WEATHER: Partly cloudy. The temperature range was 29.7oF to 66.2oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 813.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

No ash analytical reports were received during this reports coverage.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, November 28, 1998

FOR DATES: From 7:00 a.m. Friday, Nov. 27, 1998 to 7:00 a.m. Saturday, Nov. 28, 1998.

WEATHER: Overcast, clearing in the afternoon. The temperature range was 28oF to 51oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 812 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation was continued on the southern face of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-11-22-98-287) for bin 2. Bottom ash met project requirements for unrestricted backfilling on-site. Received fly ash analytical report (fly-ash-11-21-98-057) for bin A. Test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of arsenic, chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, November 27, 1998

FOR DATES: From 7:00 a.m. Thursday, Nov. 26, 1998 to 7:00 a.m. Friday, Nov. 27, 1998.

WEATHER: Rainy and overcast. The temperature range was 43oF to 54oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 14.2 hours and 480.5 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is attributed to the stoppage of hazardous waste feed while the Dry Ash Drag (DAD) was being repaired. The DAD had been stopped since 0415 hours on Thursday, November 26, 1998 due to a piece of metal debris becoming lodged within the DAD drive chain. The metal debris was removed and hazardous waste feed resumed at 1646 hours. The TDF processed contaminated soil at an average rate of 33.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance, feed preparation, lime mixing, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. No new excavation was performed in exclusion zone. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, November 26, 1998

FOR DATES: From 7:00 a.m. Wednesday, Nov. 25, 1998 to 7:00 a.m. Thursday, Nov. 26, 1998.

WEATHER: Mild with evening rain. The temperature range was 28oF to 50oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 21.2 hours and 719.2 tons of contaminated soils were thermally treated during the time frame of this report. Hazardous waste feed is currently off-line as a result of the Dry Ash Drag conveyor becoming jammed with an unknown piece of debris. The TDF processed contaminated soil at an average rate of 33.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Wastewater Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation was continued on the southern face of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program and began the weekly OSHA method 93 24-hour event for BNA. Received bottom ash analytical report (bot-ash-11-21-98-286) for bin 3. Bottom ash met project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: At 0415 hours on Thursday, November 26, 1998, hazardous waste feed to the incinerator stopped as a result of the H-510 Dry Ash Drag (DAD) conveyor locking up. It is believed that the upper portions of the DAD drive chain had become jammed with a piece of metal from the rotary kiln. OHM's thermal personnel are currently investigating the problem and plan to have the TDF back on-line by mid-morning.

V. NOTED VISITORS: Dr. Richard Stehouwer (Penn State University)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, November 25, 1998

FOR DATES: From 7:00 a.m. Tuesday, Nov. 24, 1998 to 7:00 a.m. Wednesday, Nov. 25, 1998.

WEATHER: Clear and breezy. The temperature range was 27oF to 56oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 813 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.0 tons per hour while on line. Related ongoing site activities include debris separation, feed preparation, and backfilling operations.

II. ACTIONS TAKEN: The contractor continued TDF operations, Wastewater Treatment Plant (WWTP) operations, lime mixing, debris separation, debris cleaning, wood shredding, feed preparation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were continued on the southern face of the exclusion zone. Concover, a wind and water soil erosion suppressant, was placed on the bottom ash stockpile. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash11-20-98-285) for bin 4. Test results indicated that the thermally treated soil met the project specifications for restricted backfilling on-site due to the concentration arsenic being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Denis duBreuil (USACE), Sandy Downs, Steve Webster, and Larry Newcomer (PADEP); a site tour was given to students from Bucknell University.

PREPARED BY: Steve Gibson for Bill Werntges (USACE)


Tuesday, November 24, 1998

FOR DATES: From 7:00 a.m. Monday, Nov. 23, 1998 to 7:00 a.m Tuesday, Nov. 24, 1998.

WEATHER: Warm and windy. The temperature range was 25oF to 66oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 815.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.4 tons per hour while on line. Related ongoing site activities include debris separation, feed preparation, and backfilling operations.

II. ACTIONS TAKEN: The contractor continued TDF operations, Wastewater Treatment Plant (WWTP) operations, lime mixing, debris separation, debris cleaning, feed preparation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Minor exploratory excavation was performed on the eastern side of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program. The contractor shipped a roll-off containing a composite of drum contents, previously excavated on-site, to a licensed hazardous waste treatment facility. Prior to shipment, the composite of drum contents was tested for chemical compatibility.

Received bottom ash analytical reports (bot-ash-11-18-98-283 and bot-ash-11-19-98-284) for bins 6 and 5, respectively. Test results indicated that the thermally treated soils met the project requirements for unrestricted backfilling on-site. Received fly ash analytical report (fly-ash11-18-98-056) for bin E. Test results indicated that the fly ash materials met the project specifications for restricted backfilling on-site due to the concentrations of chromium, cadmium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, November 23, 1998

FOR DATES: From 7:00 a.m. Sunday, Nov. 22, 1998 to 7:00 a.m. Monday, November 23, 1998.

WEATHER: Clear and cool. The temperature range was 26oF to 49oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 862.5 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 36.1 tons per hour while on line. Related ongoing site activities include debris separation, feed preparation, and backfill.

II. ACTIONS TAKEN: The contractor continued lime mixing, debris separation, feed preparation, TDF operations, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

Backfilled bottom ash (bot-ash-11-16-98-281) for bin 1. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site. Backfilled fly ash (flyash-11-12-98-054 and 11-15-98-055) bins C & D. The analytical test results indicated that the fly ash met project requirements for restricted backfilling above elevation 553. The concentrations of arsenic, chromium, and selenium were greater than the PADEP drinking water standard (DWS) yet less than 25 times the DWS for bin C. The concentrations of chromium and selenium were greater than the DWS yet less than 25 times the DWS for bin D.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. The contractor will be shipping the contents of drums encountered during excavation to a licensed hazardous waste handling facility for disposal. The materials were sampled and characterized prior to being manifested and shipped.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: David Modricker (USACE)


Sunday, November 22, 1998

FOR DATES: From 7:00 a.m. Saturday, Nov. 21, 1998 to 7:00 a.m. Sunday, Nov. 22, 1998.

WEATHER: Partly cloudy and breezy. The temperature range was 36oF to 41oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 860.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 36.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, and debris separation. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-11-17-98-282) for bin 7. Test results indicated that the thermally treated soil met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentration of lead greater than the PADEP drinking water standard (DWS) but less than 25 times the DWS.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, November 21, 1998

FOR DATES: From 7:00 a.m. Friday, Nov. 20, 1998 to 7:00 a.m. Saturday, Nov. 21, 1998.

WEATHER: Overcast with periods of rain. The temperature range was 37oF to 45oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.4 hours and 845.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 36.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, and debris separation. Bottom ash from bin 3 was returned to the exclusion zone during this reporting period for further thermal treatment. Excavation activities were continued on the southern face of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-11-16-98-281) for bin 1. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, November 20, 1998

FOR DATES: From 7:00 a.m. Thursday, Nov. 19, 1998 to 7:00 a.m. Friday, Nov. 20, 1998.

WEATHER: Clear and cool. The temperature range was 26oF to 50oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 863 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 36.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, waste water treatment plant (wwtp) operation, feed preparation, lime mixing, debris separation, rock cleaning, ash handling, and ash backfilling. Excavation continued along the southern face of the exclusion zone. A survey of the excavation bottom and the 2-foot rock bottom was completed on this date. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-11-15-98-280) for bin 3. Analytical tests results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site. Received fly ash analytical report (flyash-11-15-98-055) for bin D. The analytical test results indicated that the fly ash met project requirements for restricted backfilling above elevation 553 due to the concentrations of selenium and chromium being greater than the PADEP drinking water standard, but less than 25 times the standard. The target feed rate for the TDF remains at 36 tons per hour during this reporting period in an effort to optimize ash quality. It is anticipated that the feed rate will be increased in the future pending a review of operational parameters and receipt of additional ash analytical reports.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. The contractor will be shipping the contents of drums encountered during excavation to a licensed hazardous waste handling facility for disposal. The materials were sampled and characterized prior to being manifested and shipped.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: (2) Classes of Lock Haven University Students were given a presentation and a tour of the site.

PREPARED BY: David Modricker (USACE)


Thursday, November 19, 1998

FOR DATES: From 7:00 a.m. Wednesday, Nov. 18, 1998 to 7:00 a.m. Thursday, Nov. 19, 1998.

WEATHER: Clear and cool. The temperature range was 26oF to 49oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 877 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 36.7 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation continued along the southern face of the exclusion zone. A survey of the excavation bottom and the 2-foot rock bottom was completed on this date. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-11-14-98-279) for bin 3. Analytical test results indicated that the thermally treated soil did not meet the project requirements for backfilling on-site. The concentration of the semi-volatile organic compound Beta-Naphthylamine (BNA) was higher than the project criteria for backfilling. The contents of bin 3 will be returned to the exclusion zone for further thermal treatment. The target feed rate for the TDF was reduced to 36 tons per hour during this reporting period in an effort to optimize ash quality. It is anticipated that the feed rate will be increased in the future pending a review of operational parameters and receipt of additional ash analytical reports.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs and Steve Webster (PADEP)

PREPARED BY: David Modricker (USACE)


Wednesday, November 18, 1998

FOR DATES: From 7:00 a.m. Tuesday, Nov. 17, 1998 to 7:00 a.m. Wednesday, Nov. 18, 1998.

WEATHER: Overcast. The temperature range was 33oF to 50oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 948 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 39.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation continued along the southern face of the exclusion zone. Approximately 20 tons of non-RCRA, non-DOT regulated activated carbon was removed from the feed preparation building's air handling unit and shipped off-site to a carbon regeneration plant. Continued with the real time perimeter air monitoring (PAM) program. A weekly NPDES water sample was collected in the WWTP. Received bottom ash analytical report (bot-ash-11-13-98-278) for bin 4. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, November 17, 1998

FOR DATES: From 7:00 a.m. Monday, Nov. 16, 1998 to 7:00 a.m. Tuesday, Nov. 17, 1998.

WEATHER: Overcast. The temperature range was 32oF to 48oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 949.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 39.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation continued along the southern face of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-11-12-98-277) for bin 5. Test results indicated that the thermally treated soil met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentration of arsenic being greater than the PADEP drinking water standard, but less than 25 times the standard. Received fly ash analytical report (fly-ash-11-12-98-054) for bin C. Test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentration of arsenic, selenium, and chromium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, November 16, 1998

FOR DATES: From 7:00 a.m. Sunday, Nov. 15, 1998 to 7:00 a.m. Monday, Nov. 16, 1998.

WEATHER: Mostly sunny. The temperature range was 35oF to 55oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 951.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris separation, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program and recovered OSHA 93 samples and NIOSH 5518 samples for BNA.

Received bottom ash analytical reports (bot-ash-11-10-98-275 & bot-ash-11-11-98-276). Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for David Modricker (USACE)


Sunday, November 15, 1998

FOR DATES: From 7:00 a.m. Saturday, Nov. 14, 1998 to 7:00 a.m. Sunday, Nov. 15, 1998.

WEATHER: Overcast and cool with a trace of rain. The temperature range was 31oF to 57oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 950.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 39.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris separation, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program and initiated OSHA 93 sampling and NIOSH 5518 sampling for BNA at both the on-site and off-site sampling locations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, November 14, 1998

FOR DATES: From 7:00 a.m. Friday, Nov. 13, 1998 to 7:00 a.m. Saturday, Nov. 14, 1998.

WEATHER: Overcast and cool. The temperature range was 30oF to 50oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 952.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. Excavation was continued on the southern face of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program and continued with the audit of the air monitoring systems.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, November 13, 1998

FOR DATES: From 7:00 a.m. Thursday, Nov. 12, 1998 to 7:00 a.m. Friday, Nov. 13, 1998.

WEATHER: Clear, cool and breezy. The temperature range was 31oF to 52oF.

I. CURRENT SITUATION: Contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 957.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation

II. ACTIONS TAKEN: The Contractor continued TDF operations, Wastewater Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. Excavation activities continued along the southern face of the exclusion zone. The contractor continued with the real time Perimeter Air Monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-11-8-98-274) for bin 1. Test results indicated that the thermally treated soil met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentration of lead being greater than the PADEP drinking water standard, but less than 25 times the standard. Received fly ash analytical report (fly-ash-11-8-98-053) for bin B. Test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP), Dr. S. Tucker & R. Lion (IT Corporation)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Thursday, November 12, 1998

FOR DATES: From 7:00 a.m. Wednesday, Nov. 11, 1998 to 7:00 a.m. Thursday, Nov. 12, 1998.

WEATHER: Clear, cool and windy. The temperature range was 34oF to 55oF.

I. CURRENT SITUATION: Contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 914.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 38.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation

II. ACTIONS TAKEN: The Contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation continued along the southern face of the exclusion zone. Sheetpiling along the Conrail side of the site was cut to grade elevation. The contractor continued with the real time Perimeter Air Monitoring (PAM) program. The air monitoring subcontractor, MRI, began conducting their annual quality assurance audit of the off-site and perimeter air monitoring systems.

Received bottom ash analytical report (bot-ash-11-7-98-273) for bin 2. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: J. Cheney (USACE), Dr. S. Tucker & R. Lion (IT), Jack Balsinger & Jeff Kittel (MRI).

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, November 11, 1998

FOR DATES: From 7:00 a.m. Tuesday, Nov. 10, 1998 to 7:00 a.m. Wednesday, Nov. 11, 1998.

WEATHER: Overcast, cool, wind gusts & inter. showers. The temperature range was 39oF to 61oF.

I. CURRENT SITUATION: Contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). Hazardous waste feed to the incinerator resumed at 1134 hours on Tuesday, November 10, 1998. The TDF was on-line for 19.3 hours and 763.3 tons of contaminated soils were thermally treated. The TDF processed contaminated soil at an average rate of 39.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation

II. ACTIONS TAKEN: The Contractor continued TDF operations, WWTP operations, feed preparation, excavation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor continued with the real time Perimeter Air Monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-11-6-98-272) for bin 3. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547.

MRI, OHM, & USACE participated in Off-site Air Monitoring sampling QA observations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: J. Cheney-USACE, S. Tucker- IT, Jack Balsinger & Jeff Kittel-MRI, Sandra Downs - PADEP.

PREPARED BY: Danny Zion for David Modricker (USACE)


Tuesday, November 10, 1998

FOR DATES: From 7:00 a.m. Monday, Nov. 9, 1998 to 7:00 a.m. Tuesday, Nov. 10, 1998.

WEATHER: Overcast and cool with evening showers. The temperature range was 31oF to 48oF.

I. CURRENT SITUATION: Hazardous waste feed to the incinerator was stopped at 0822 hours on Monday, November 9, 1998 in order to perform monthly interlock checks, maintenance to the dual feed screws and make repairs to the dry ash drag conveyor. Prior to stopping hazardous waste feed to the incinerator, the Thermal Destruction Facility (TDF) was on-line for 1.4 hours and 54.3 tons of contaminated soils were thermally treated. The TDF processed contaminated soil at an average rate of 38.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation. Hazardous waste feed to the incinerator is tentatively scheduled to resume during the morning hours on Tuesday, November 10, 1998.

II. ACTIONS TAKEN: The TDF was placed in a 'hot hold' mode to facilitate interlock checks, weld carbide cutting teeth on the existing feed screw flights and repair wear damage to the floor plate of the Dry Ash Drag. Interlock checks are performed monthly in order to test the TDF's automatic shut down systems. As part of the interlock checks, a planned Thermal Relief Valve (TRV) opening was initiated at 1404 hours. Prior to the TRV opening, feed to the incinerator had been stopped for approximately 5.5 hours. Proper regulatory authorities were notified by OHM. The contractor continued WWTP operations, feed preparation, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Minor excavation was performed in the exclusion zone. The contractor continued with the real time Perimeter Air Monitoring (PAM) program and finished weekly OSHA method 93 tests for BNA at each PAM station and the off-site air monitoring stations.

Received bottom ash analytical report (bot-ash-11-5-98-271) for bin 4. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical report (fly-ash-11-4-98-052) for bin A. Test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the levels of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, November 9, 1998

FOR DATES: From 7:00 a.m. Sunday, Nov. 8, 1998 to 7:00 a.m. Monday, Nov. 9, 1998.

WEATHER: Overcast and cool with some rain showers. The temperature range was 40oF to 46oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.6 hours and 942.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, ash handling, and ash backfilling.

Received bottom ash analytical report (bot-ash-11-3-98-269 and bot-ash-11-4-98-270). Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). No fly ash analytical reports were received during this reporting period.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a TDF shutdown (hot hold) tentatively scheduled for November 9, 1998 in order to perform monthly interlock checks and perform maintenance to the existing dual feed screws. A planned thermal relief valve opening is scheduled for Nov 9, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for David Modricker (USACE)


Sunday, November 8, 1998

FOR DATES: From 7:00 a.m. Saturday, Nov. 7, 1998 to 7:00 a.m. Sunday, Nov. 8, 1998.

WEATHER: Overcast and cool with some rain showers. The temperature range was 40oF to 48oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 957.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, rock cleaning, ash handling, and ash backfilling. Continued with real time perimeter air monitoring (PAM). Work continues to repair the heat trace line to PAM station 3. Backup sampling is conducted when the primary sampling system is under repair.

Received bottom ash analytical report (bot-ash-11-2-98-268) for bin 7. Test results indicated that the thermally treated soil met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the levels of lead being greater than the PADEP drinking water standard, but less than 25 times the standard. No fly ash analytical reports were received during this reports coverage.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a TDF shutdown (hot hold) tentatively scheduled for November 9, 1998 in order to perform monthly interlock checks and perform maintenance to the existing dual feed screws. A planned thermal relief valve opening is scheduled for Nov 9, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, November 7, 1998

FOR DATES: From 7:00 a.m. Friday, Nov. 6, 1998 to 7:00 a.m. Saturday, Nov. 7, 1998.

WEATHER: Overcast cloudy and cool. The temperature range was 38oF to 43oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 959.0 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation resumed on the southern end of the exclusion zone. Continued with real time perimeter air monitoring (PAM). Work continues to repair the heat trace line to PAM station 3. Backup sampling is conducted when the primary sampling system is under repair.

Received bottom ash analytical report (bot-ash-11-1-98-267) for bin 1. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site. Received fly ash analytical report (fly-ash-11-1-98-051) for bin E. Test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the levels of arsenic, chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a TDF shutdown (hot hold) tentatively scheduled for November 9, 1998 in order to perform monthly interlock checks and perform maintenance to the existing dual feed screws. A planned thermal relief valve opening is scheduled for Nov 9, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, November 6, 1998

FOR DATES: From 7:00 a.m. Thursday, Nov. 5, 1998 to 7:00 a.m. Friday, Nov. 6, 1998.

WEATHER: Overcast and cool. The temperature range was 34oF to 44oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 959.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. Minor exploratory excavation was performed on this date. The contractor continued to erect the lime silo on the ash stabilization pad within the exclusion zone. Continued with the real time Perimeter Air Monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a TDF shutdown (hot hold) tentatively scheduled for November 9, 1998 in order to perform monthly interlock checks and perform maintenance to the existing dual feed screws. A planned thermal relief valve opening is scheduled for Nov 9, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for William Werntges (USACE)


Thursday, November 5, 1998

FOR DATES: From 7:00 a.m. Wednesday, Nov. 4, 1998 to 7:00 a.m. Thursday, Nov. 5, 1998.

WEATHER: Clear and cool. The temperature range was 21oF to 43oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 957.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. A survey of the 2-foot rock bottom was performed. No new excavation was performed on this date, rather the existing feed stockpile was graded to original ground. The contractor continued to erect the lime silo on the ash stabilization pad within the exclusion zone. The primary Perimeter Air Monitoring (PAM) system was brought back on-line. Received bottom ash analytical report (bot-ash-10-31-98-266) for bin 2. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a TDF shutdown (hot hold) tentatively scheduled for November 9, 1998 in order to perform monthly interlock checks and perform maintenance to the existing dual feed screws. A planned thermal relief valve opening is scheduled for Nov 9, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, November 4, 1998

FOR DATES: From 7:00 a.m. Tuesday, Nov. 3, 1998 to 7:00 a.m. Wednesday, Nov. 4, 1998.

WEATHER: Clear. The temperature range was 21oF to 49oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 955.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. No new excavation was performed on this date. The contractor erected the lime silo on the ash stabilization pad within the exclusion zone. The primary Perimeter Air Monitoring (PAM) system remained off-line in order to investigate and make repairs to the heat trace line connected to PAM Station 3. During the repair period, perimeter site air was monitored with hand held devices. The monthly on-site test for TSPs and metals was initiated at the PAM stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a TDF shutdown (hot hold) tentatively scheduled for November 9, 1998 in order to perform monthly interlock checks and perform maintenance to the existing dual feed screws. A planned thermal relief valve opening is scheduled for Nov 9, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs and Steve Webster (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, November 3, 1998

FOR DATES: From 7:00 a.m. Monday, Nov. 2, 1998 to 7:00 a.m. Tuesday, Nov. 3, 1998.

WEATHER: Partly cloudy. The temperature range was 25oF to 53oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 952.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. No new excavation was performed on this date. The contractor performed a survey of the excavation bottom. The primary Perimeter Air Monitoring (PAM) system was taken off-line in order to investigate and make repairs to the heat trace line connected to PAM Station 3. During the repair period, perimeter site air was monitored with hand held devices. The contractor began the weekly OSHA method 93 24-hour air sampling event for BNA at the off-site air and perimeter air monitoring stations.

Received bottom ash analytical reports (bot-ash-10-27-98-262, bot-ash-10-28-98-263 and bot-ash-10-29-98-264) for bin 6, bin 5 and bin 4 respectively. Analytical test results indicate that the thermally treated soil met the project requirements for unrestricted backfilling on-site. Received fly ash analytical report (fly-ash-10-28-98-050) for bin D. Test results indicated that the fly ash material met the project requirements for restricted backfilling (i.e. above elevation 553) due to the concentrations of selenium and chromium being greater than the PADEP drinking water standards, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a TDF shutdown (hot hold) tentatively scheduled for November 9, 1998 in order to perform monthly interlock checks and perform maintenance to the existing dual feed screws. The installation of a lime silo storage unit will begin this week.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, November 2, 1998

FOR DATES: From 7:00 a.m. Sunday, Nov. 1, 1998 to 7:00 a.m. Monday, Nov. 2, 1998.

WEATHER: Partly sunny and mild. The temperature range was 29oF to 57oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 955.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 39.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. No new excavation was performed on this date. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. The installation of a lime silo storage unit will begin this week. It is expected that repairs to a heat traced intake line of the PAM, as well as maintenance on the primary system will occur today. The backup system will be used while the primary PAM system is repaired.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Chad Thompson for David Modricker (USACE)


Sunday, November 1, 1998

FOR DATES: From 7:00 a.m. Saturday, Oct. 31, 1998 to 7:00 a.m. Sunday, Nov. 1, 1998.

WEATHER: Mostly sunny and breezy. The temperature range was 270F to 620F.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 958.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard L. Lupold for David Modricker (USACE)

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.